P varies inversely as h. If P = 7 when h = 0.7, determine P when h = 12.
Answer by jim_thompson5910(35256)

"P varies inversely as h" ----> P = k/h

P = k/h

7 = k/0.7

7*0.7 = k

4.9 = k

k = 4.9

-------------

P = k/h

P = 4.9/h

P = 4.9/12

P = 0.408333333 (this is approximate, the 3's repeat forever)

Note: 49/120 = 0.408333333 (approximately, the 3's repeat forever)

So you can report the answer as P = 49/120